Eliminate the arbitrary constants indicated in brackets from the following equation

and form corresponding partial differential equation:


e

1

{z−(

x2

y

)}

=

ax

2

y2 +

b

y

;(a,b)


1
Expert's answer
2021-01-17T13:56:05-0500

"\displaystyle
e^{z - x^2y} = ax^2y^2 + by\\

e^{z - x^2y}\left(\frac{\partial z}{\partial x} - 2xy\right) = 2axy^2\\

e^{z - x^2y}\left(\frac{\partial z}{\partial y} - x^2\right) = 2ax^2y + b\\

e^{z - x^2y}\left(\left(\frac{\partial z}{\partial x} - 2xy\right)^2 + \frac{\partial^2 z}{\partial x^2} - 2x\right) = 2ay^2\\


e^{z - x^2y}\left(\left(\frac{\partial z}{\partial y} - x^2\right)^2 + \frac{\partial^2 z}{\partial y^2}\right) = 2ax^2\\

\frac{e^{z - x^2y}}{y^2}\left(\left(\frac{\partial z}{\partial x} - 2xy\right)^2 + \frac{\partial^2 z}{\partial x^2} - 2x\right) = \frac{e^{z - x^2y}}{x^2}\left(\left(\frac{\partial z}{\partial y} - x^2\right)^2 + \frac{\partial^2 z}{\partial y^2}\right)\\


x^2\left(\frac{\partial z}{\partial x} - 2xy\right)^2 + x^2\frac{\partial^2 z}{\partial x^2} - 2x^3 = y^2\left(\frac{\partial z}{\partial y} - x^2\right)^2 + y^2\frac{\partial^2 z}{\partial y^2}\\

\left(x\frac{\partial z}{\partial x} - 2x^2y - y\frac{\partial z}{\partial y} + x^2y\right)\left(x\frac{\partial z}{\partial x} - 2x^2y + y\frac{\partial z}{\partial y} - x^2y\right) + x^2\frac{\partial^2 z}{\partial x^2} - y^2\frac{\partial^2 z}{\partial y^2} - 2x^3 = 0"